112 年 國立政治大學金融學系碩士班金融管理組《統計學A》

📄 試題原卷 免費註冊後即可對照原始考卷 PDF免費註冊

第 1 題25 分

The table below is the total number of phone calls arriving to the information center of a bank in 20 randomly chosen working days during its open period (9:00-17:00).

Time9:00:00-9:59:5910:00:00-10:59:5911:00:00-11:59:5912:00:00-12:59:5913:00:00-13:59:5914:00:00-14:59:5915:00:00-15:59:5916:00:00-16:59:59
Total number of calls170285314370299122194236

The calls occur randomly and independently of one another.

(1) Please estimate the average number of calls in an hour. (5%)
(2) Assume that the hourly number of calls arriving to this information center follows a Poisson distribution. Find the probability that the information center receives more than two calls (including) at 11:00:00-11:10:00 (10 minutes) on a working day. (10%)
(3) Let pp represent the ratio of the number of calls arriving to this information center at 12:00:00-12:59:59 to the total number of calls in all working days. At the significance level of 0.01, test whether the proportion of calls, pp, exceeds 15%. (10%)

登入後即可作答並保存紀錄。

這一題的完整詳解

本題主要考驗對平均數的估計、Poisson 分布的應用以及比例檢定。

(1) 估計每小時的平均來電數

首先,我們需要計算這 20 個工作日總共的來電數。
總來電數 =170+285+314+370+299+122+194+236=1990= 170 + 285 + 314 + 370 + 299 + 122 + 194 + 236 = 1990 通。

這些來電發生在 20 個工作日,每個工作日的時間長度為 9:00 到 17:00,共 8 小時。
因此,總觀察時間為 20 天×8 小時/天=160 小時20 \text{ 天} \times 8 \text{ 小時/天} = 160 \text{ 小時}。

估計的平均每小時來電數為總來電數除以總觀察小時數:
平均每小時來電數=1990160=12.4375\text{平均每小時來電數} = \frac{1990}{160} = 12.4375 通/小時。

【答案】12.4375

(2) Poisson 分布與機率計算

題目假設每小時的來電數服從 Poisson 分布。
由第一小題,我們估計的平均每小時來電數為 λhour=12.4375\lambda_{hour} = 12.4375。
Poisson 分布的參數 λ\lambda 代表在一個給定的時間間隔內事件發生的平均次數。

我們需要計算在 11:00:00 到 11:10:00 這 10 分鐘內,收到超過兩通電話的機率。
這 10 分鐘是 1 小時的 1060=16\frac{10}{60} = \frac{1}{6}。
因此,在這個 10 分鐘的時間間隔內,平均來電數 λ10min\lambda_{10min} 為:
λ10min=λhour×1060=12.4375×16≈2.0729\lambda_{10min} = \lambda_{hour} \times \frac{10}{60} = 12.4375 \times \frac{1}{6} \approx 2.0729。

設 XX 為在 10 分鐘內收到的電話次數,X∼Poisson(λ=2.0729)X \sim \text{Poisson}(\lambda = 2.0729)。
我們要求 P(X>2)P(X > 2)。
P(X>2)=1−P(X≤2)=1−[P(X=0)+P(X=1)+P(X=2)]P(X > 2) = 1 - P(X \le 2) = 1 - [P(X=0) + P(X=1) + P(X=2)]。

Poisson 機率公式為 P(X=k)=e−λλkk!P(X=k) = \frac{e^{-\lambda} \lambda^k}{k!}。

P(X=0)=e−2.0729(2.0729)00!=e−2.0729≈0.1259P(X=0) = \frac{e^{-2.0729} (2.0729)^0}{0!} = e^{-2.0729} \approx 0.1259
P(X=1)=e−2.0729(2.0729)11!=2.0729×e−2.0729≈2.0729×0.1259≈0.2611P(X=1) = \frac{e^{-2.0729} (2.0729)^1}{1!} = 2.0729 \times e^{-2.0729} \approx 2.0729 \times 0.1259 \approx 0.2611

🔒

後續完整解題步驟與【答案】

免費註冊,享三天全站完整詳解閱覽。

免費註冊

第 2 題20 分

A quality control engineer is interested in the mean length of sheet insulation being cut automatically by machine. The desired length of the sheet insulation is 12 feet.

(1) The engineer randomly chooses 20 sheet insulations and the resulting sample mean length is 12.15 and sample standard deviation in the cutting length is 0.2 feet. Test whether the cutting machine is out of control at the significance level of 0.05. (10%)
(2) Suppose the engineer decided to estimate the mean length to within 0.025 with 99% confidence. What sample size would be needed? (10%)

登入後即可作答並保存紀錄。

這一題的完整詳解

本題主要考驗對假設檢定(t檢定)與樣本大小的估計。

(1) 檢定機器是否失控(t 檢定)

題目要求檢定機器的平均切割長度是否偏離目標值 12 英尺。
我們有樣本資料:
樣本數 n=20n = 20
樣本平均數 xˉ=12.15\bar{x} = 12.15 英尺
樣本標準差 s=0.2s = 0.2 英尺
目標值 μ0=12\mu_0 = 12 英尺
顯著水準 α=0.05\alpha = 0.05

由於母體標準差未知,且樣本數 n=20n=20 較小,我們使用 t 檢定。
虛無假設 H0:μ=12H_0: \mu = 12 (機器在控制中,平均長度為 12 英尺)
對立假設 H1:μ≠12H_1: \mu \ne 12 (機器失控,平均長度不為 12 英尺)

這是一個雙尾檢定。
檢定統計量為 t=xˉ−μ0s/nt = \frac{\bar{x} - \mu_0}{s/\sqrt{n}}。
t=12.15−120.2/20=0.150.2/20=0.150.2/4.472=0.150.04472≈3.354t = \frac{12.15 - 12}{0.2/\sqrt{20}} = \frac{0.15}{0.2/\sqrt{20}} = \frac{0.15}{0.2/4.472} = \frac{0.15}{0.04472} \approx 3.354。

自由度 df=n−1=20−1=19df = n - 1 = 20 - 1 = 19。
在顯著水準 α=0.05\alpha = 0.05 下,雙尾檢定的臨界值 tα/2,dft_{\alpha/2, df}。
我們查詢 t 分布表,對於 df=19df=19 和 α/2=0.025\alpha/2 = 0.025,臨界值為 t0.025,19≈2.093t_{0.025, 19} \approx 2.093。

比較檢定統計量與臨界值:
∣t∣=∣3.354∣≈3.354|t| = |3.354| \approx 3.354。
因為 ∣t∣≈3.354>tcritical≈2.093|t| \approx 3.354 > t_{critical} \approx 2.093,我們拒絕虛無假設 H0H_0。
結論是,有足夠的證據顯示該切割機器的平均切割長度不等於 12 英尺,即機器失控。

【答案】拒絕虛無假設,證明機器失控。

(2) 估計平均長度所需的樣本大小

🔒

後續完整解題步驟與【答案】

免費註冊,享三天全站完整詳解閱覽。

免費註冊

第 3 題27 分

The following table classifies students by field of study and whether or not they have a loan. The students are randomly selected from a large university.

Field of studyYesNo
Education2036
Engineering94141
Management2451
Science5139
Liberal arts79125

(1) At the significance level of 0.05, test whether the proportion of students studying in Engineering having a loan is greater than that of students studying in Management having a loan. (12%)
(2) Conduct a test to examine whether all the proportions of students having a loan in different fields are equal at the significance level of 0.1. (15%)

登入後即可作答並保存紀錄。

這一題的完整詳解

本題主要考驗對兩個獨立樣本比例檢定以及卡方獨立性檢定(或多個比例檢定)。

首先,我們計算各欄位的總人數:

Field of studyYes (Loan)No (No Loan)Total
Education203656
Engineering94141235
Management245175
Science513990
Liberal arts79125204
Total268392760

(1) 檢定工程系學生有貸款比例是否大於管理系學生有貸款比例

此為檢定兩個獨立比例的檢定。
我們定義:
pEp_E: 工程系學生有貸款的比例
pMp_M: 管理系學生有貸款的比例

樣本資料:
工程系 (E):樣本數 nE=235n_E = 235,有貸款人數 xE=94x_E = 94。樣本比例 p^E=94235≈0.4000\hat{p}_E = \frac{94}{235} \approx 0.4000。
管理系 (M):樣本數 nM=75n_M = 75,有貸款人數 xM=24x_M = 24。樣本比例 p^M=2475=0.3200\hat{p}_M = \frac{24}{75} = 0.3200。

顯著水準 α=0.05\alpha = 0.05。
虛無假設 H0:pE≤pMH_0: p_E \le p_M (或 pE−pM≤0p_E - p_M \le 0)
對立假設 H1:pE>pMH_1: p_E > p_M (或 pE−pM>0p_E - p_M > 0)

這是右尾檢定。
我們需要計算合併比例 p^pool=xE+xMnE+nM=94+24235+75=118310≈0.3806\hat{p}_{pool} = \frac{x_E + x_M}{n_E + n_M} = \frac{94 + 24}{235 + 75} = \frac{118}{310} \approx 0.3806。

檢定統計量為 z=p^E−p^Mp^pool(1−p^pool)(1nE+1nM)z = \frac{\hat{p}_E - \hat{p}_M}{\sqrt{\hat{p}_{pool}(1-\hat{p}_{pool})(\frac{1}{n_E} + \frac{1}{n_M})}}。
z=0.4000−0.32000.3806(1−0.3806)(1235+175)z = \frac{0.4000 - 0.3200}{\sqrt{0.3806(1-0.3806)(\frac{1}{235} + \frac{1}{75})}}
z=0.08000.3806×0.6194×(1235+175)z = \frac{0.0800}{\sqrt{0.3806 \times 0.6194 \times (\frac{1}{235} + \frac{1}{75})}}
z=0.08000.2358×(0.004255+0.013333)z = \frac{0.0800}{\sqrt{0.2358 \times (0.004255 + 0.013333)}}
z=0.08000.2358×0.017588z = \frac{0.0800}{\sqrt{0.2358 \times 0.017588}}
z=0.08000.004150z = \frac{0.0800}{\sqrt{0.004150}}
z=0.08000.06442≈1.2418z = \frac{0.0800}{0.06442} \approx 1.2418。

在顯著水準 α=0.05\alpha = 0.05 下,右尾檢定的臨界值 zcriticalz_{critical}。
z0.05≈1.645z_{0.05} \approx 1.645。

比較檢定統計量與臨界值:
z≈1.2418<zcritical=1.645z \approx 1.2418 < z_{critical} = 1.645。
我們不拒絕虛無假設 H0H_0。
結論是,沒有足夠的證據顯示工程系學生有貸款的比例大於管理系學生。

【答案】不拒絕虛無假設。

(2) 檢定不同科系有貸款比例是否相等(卡方獨立性檢定)

此題是檢定一個類別變數(Field of study)與另一個類別變數(Student loan: Yes/No)之間是否獨立。若兩者獨立,則各科系有貸款的比例是相同的。
我們使用卡方獨立性檢定。
顯著水準 α=0.1\alpha = 0.1。

虛無假設 H0H_0: 所有科系有貸款的比例都相等。
對立假設 H1H_1: 至少有一個科系有貸款的比例不等於其他科系。

首先,我們計算期望次數 (Expected counts)。
總樣本數 N=760N = 760。
總有貸款人數 NYes=268N_{Yes} = 268。
總無貸款人數 NNo=392N_{No} = 392。

合併比例 p^pool=NYesN=268760≈0.3526\hat{p}_{pool} = \frac{N_{Yes}}{N} = \frac{268}{760} \approx 0.3526。
合併無貸款比例 p^pool,No=NNoN=392760≈0.6474\hat{p}_{pool, No} = \frac{N_{No}}{N} = \frac{392}{760} \approx 0.6474。

預期次數 Eij=(列總計)×(行總計)NE_{ij} = \frac{(\text{列總計}) \times (\text{行總計})}{N}。

| Field of study | Yes (Observed) | No (Observed) | Total | Yes (Expected) | No (Expected) |

🔒

後續完整解題步驟與【答案】

免費註冊,享三天全站完整詳解閱覽。

免費註冊

第 4 題28 分

A bank manager considers a two-factor experimental design to examine whether different age groups (Age with 4 levels) of customers and interface designs (Interface with 3 levels) of the automated teller machine (ATM) have effect on the operation time (in seconds) when their customers use an ATM. Each treatment has 3 randomly chosen customers. The manager obtains the following analysis of variance (ANOVA) table.

Sum of squaresDegrees of freedomMean squaresF
Age64.4(B)(G)
Interface27.6(C)(H)
Age × Interface135.4(D)(I)
Error(A)(E)(J)
Total399.5(F)

(1) Compute the values of (A) to (M), and describe their required formulas and calculation process in details. (18%)
(2) Describe the null hypothesis and alternative hypothesis for (L) and (M). Draw your conclusions for both tests at a significant level of 0.05. (10%)

For all questions about statistical tests, please specify null hypothesis and alternative hypothesis in details.

登入後即可作答並保存紀錄。

這一題的完整詳解

本題主要考驗對 ANOVA 表格的理解、計算以及相關假設檢定的設定與解釋。

(1) 計算 ANOVA 表格中的值 (A) 到 (M)

ANOVA 表格中的各項值是通過以下關係計算的:

  • Sum of Squares (SS): 各項效應的變異量。
  • Degrees of Freedom (df): 各項效應的自由度。
  • Mean Squares (MS): 變異量除以自由度,即 MS=SS/dfMS = SS / df。
  • F-statistic: 不同效應的 MS 與 Error MS 的比值,即 F=MSeffect/MSerrorF = MS_{effect} / MS_{error}。

我們已知:

  • Age SS = 64.4
  • Interface SS = 27.6
  • Age × Interface SS = 135.4
  • Total SS = 399.5
  • Age levels = 4
  • Interface levels = 3
  • Number of customers per treatment = 3

首先,計算各項的自由度 (df):

  • Age df (B): 因子 Age 的水準數減 1。 dfAge=(Age levels)−1=4−1=3df_{Age} = (\text{Age levels}) - 1 = 4 - 1 = 3。
  • Interface df (C): 因子 Interface 的水準數減 1。 dfInterface=(Interface levels)−1=3−1=2df_{Interface} = (\text{Interface levels}) - 1 = 3 - 1 = 2。
  • Age × Interface df (D): 兩個因子 df 的乘積。 dfAge×Interface=dfAge×dfInterface=3×2=6df_{Age \times Interface} = df_{Age} \times df_{Interface} = 3 \times 2 = 6。
  • Total df (F): 總樣本數減 1。總樣本數 = (Age levels) ×\times (Interface levels) ×\times (Customers per treatment) =4×3×3=36= 4 \times 3 \times 3 = 36。
    dfTotal=(Total samples)−1=36−1=35df_{Total} = (\text{Total samples}) - 1 = 36 - 1 = 35。
  • Error df (E): 總自由度減去其他效應的自由度。 dfError=dfTotal−dfAge−dfInterface−dfAge×Interfacedf_{Error} = df_{Total} - df_{Age} - df_{Interface} - df_{Age \times Interface}。
    dfError=35−3−2−6=24df_{Error} = 35 - 3 - 2 - 6 = 24。
  • (A): Error SS。總 SS 等於所有效應 SS 的總和。
    SSTotal=SSAge+SSInterface+SSAge×Interface+SSErrorSS_{Total} = SS_{Age} + SS_{Interface} + SS_{Age \times Interface} + SS_{Error}。
    399.5=64.4+27.6+135.4+SSError399.5 = 64.4 + 27.6 + 135.4 + SS_{Error}。
    SSError=399.5−(64.4+27.6+135.4)=399.5−227.4=172.1SS_{Error} = 399.5 - (64.4 + 27.6 + 135.4) = 399.5 - 227.4 = 172.1。
    所以 (A) = 172.1。

接下來,計算各項的 Mean Squares (MS):

  • Age MS (G): MSAge=SSAge/dfAge=64.4/3≈21.467MS_{Age} = SS_{Age} / df_{Age} = 64.4 / 3 \approx 21.467。
  • Interface MS (H): MSInterface=SSInterface/dfInterface=27.6/2=13.8MS_{Interface} = SS_{Interface} / df_{Interface} = 27.6 / 2 = 13.8。
  • Age × Interface MS (I): MSAge×Interface=SSAge×Interface/dfAge×Interface=135.4/6≈22.567MS_{Age \times Interface} = SS_{Age \times Interface} / df_{Age \times Interface} = 135.4 / 6 \approx 22.567。
  • Error MS (J): MSError=SSError/dfError=172.1/24≈7.171MS_{Error} = SS_{Error} / df_{Error} = 172.1 / 24 \approx 7.171。

最後,計算 F 統計量:

  • Age F (K): FAge=MSAge/MSError=21.467/7.171≈2.993F_{Age} = MS_{Age} / MS_{Error} = 21.467 / 7.171 \approx 2.993。
🔒

後續完整解題步驟與【答案】

免費註冊,享三天全站完整詳解閱覽。

免費註冊

其他考古題